# Model-free Forecasting of Rogue Waves using Reservoir Computing

### v1

@@ 1 Introduction
Reservoir Computing (RC) is a specific variant of the recurrent neural network (RNN) in machine learning [22]. RC can be viewed as a three-layer RNN: an input layer, an internal layer state (the reservoir), and an output layer. In contrast to the conventional RNN, the reservoir parameters are fixed, sparse, and random. Surprisingly, despite having a simplified learning procedure compared to RNNs, RC has been reported to be more powerful [23, 24, 25]. The framework has successfully modeled both short-term and long-term dynamics of chaotic systems, capturing their attractors [26, 27, 28, 29, 30]. An extension of RC for learning spatiotemporal problems was presented in [27, 25, 31, 32, 33, 34, 35]. Recent studies [36, 37] have introduced knowledge-based models within the RC framework. These advances highlight the versatility and potential of RC in addressing a wide range of spatiotemporal challenges. Another motivation for RC lies in its potential for physical implementation, as reviewed in [38, 39, 40], which has the potential to achieve both computational speedup and improved energy efficiency.
This research investigates the prediction and forecasting capabilities of RC to predict short-term nonlinear dynamics of integrable systems modeled by the NLS equation. Despite various studies on RC in chaotic spatiotemporal systems, little attention has been given to its application to Hamiltonian counterparts. An investigation into RC’s ability to replicate the phase plane of Hamiltonian systems is reported in [41], although this study is limited to low-dimensional systems. Only [32] attempted to simulate NLS using RC, focusing on periodic Akhmadiev breathers, Kutnezov-Ma solitons, and second-order periodic breather collisions. Despite this insight, the similarity between the training and testing data in [32] raises questions about the model’s generalizability.
@@
The primary objective of this research is to advance the understanding of RC’s forecasting capabilities for non-periodic waves, particularly rogue waves. A key focus lies in evaluating its performance using testing data from a different initial condition, aiming to thoroughly assess its robustness and applicability in varying scenarios. While several studies have demonstrated RC’s ability to predict unseen dynamics in chaotic dissipative systems [42, 29, 43], its performance in other systems—particularly Hamiltonian systems—remains largely unexplored due to the absence of an attractor. Our main contribution to the project is the revelation that RC can provide reasonable predictions, even when operating within different phase spaces in the integrable spatiotemporal system. Furthermore, we introduce a modified forecasting approach that incorporates real data assimilation, significantly improving RC’s predictive performance. This result paves the way for future research to better understand and improve the long-term predictive capabilities of RC within Hamiltonian systems.
This report is organized as follows. Section 2 presents a brief introduction to parallel RC and details the training and testing procedures used in the study. In Section 3, we provide a detailed explanation of the training data and the prediction results. The main focus of our study is explained in Sections 4 and 5. In these sections, we employ different testing data to assess the generalizability of the RC model trained in the previous section for higher-order breathers and random ocean waves. Section 6 discusses procedures to extend the prediction horizon in autonomous prediction. We conclude in Section 7 and provide an overview of future research directions.
2 RC Framework
@@ 2.1 Parallel RC
We will refer to the pair $(f,g)$ as an RC, where the input states are denoted as $u_{j}\in\mathbb{R}^{d_{I}}$ and the hidden states as $x_{j}\in\mathbb{R}^{d_{S}}$. As evident from Eqs. (1)-(2), RC is driven by a sequence of input states $\mathbf{u}=(\ldots,u_{-1},u_{0})$, complemented by an initial hidden state $x_{-\infty}$. We denote $\mathbf{x}=(\ldots,x_{-1},x_{0})$ and write $\mathbf{x}=f(\mathbf{u},x_{-\infty})$ if each element of the sequence $\mathbf{x}$ is obtained by applying the operation in (1) to the input sequence $\mathbf{u}$. Consequently, the output sequence is written as $\mathbf{y}=(f,g)(\mathbf{u})$.
When the effect of the initial hidden state $x_{-\infty}$ becomes negligible after a certain number of iterations, implying that the hidden state depends solely on the input sequence, the dynamical system exhibits the Echo State Property (ESP) [23]. Technically, if $\mathbf{x}=f(\mathbf{u},x_{-\infty})$ and $\mathbf{\tilde{x}}=f(\mathbf{u},\tilde{x}_{-\infty})$ are two sequences of hidden states corresponding to different initial conditions, the ESP condition implies that $x_{0}=\tilde{x}_{0}$.
@@
The choice of RC, as described by Eqs. (3)-(4), takes into account the symmetry inherent in the underlying data to be simulated. Specifically, the RC exhibits symmetry such that $(f,g)(\mathbf{u})=-(f,g)(-\mathbf{u})$. The importance of preserving the symmetry of the RC output, which mirrors the symmetry of the original system, has been emphasized in previous works [27, 45]. Acknowledging and preserving this symmetry is crucial for ensuring the fidelity of the simulated dynamics and improving the ability of the RC model to accurately capture the underlying patterns and behaviors of the system being studied.
@@ 2.2 Training Procedure
Figure 1: Illustration of the parallel RC architecture. The spatial grid is divided into overlapping grids $u^{(k)}$ as input to the RCs and non-overlapping grids $y^{(k)}$ as the output.
@@
where $N_{w}$ is the number of initialization steps used to ”warm up” the RCs. The first $N_{w}$ hidden states are discarded to eliminate any dependencies on initialization. The terms $\hat{y}_{j}^{(k)}$ represent the true output data, and $\beta$ is the Tikhonov parameter that controls the regularization strength of the least-squares problem. The norm of the matrix $W_{o}$ is computed using the Frobenius norm.
@@
Let us define $\mathbf{X}^{(k)}$ and $\mathbf{\hat{Y}}^{(k)}$ as matrices whose columns are the vectors $x_{j}^{(k)}$ and $\hat{y}_{j}^{(k)}$, respectively, for $-N+N_{w}\leq j\leq 0$. Furthermore, we construct the matrix $\mathbf{X}$ by stacking the matrices $\mathbf{X}^{(k)}$ column-wise, and similarly for $\mathbf{\hat{Y}}$. The loss function in Eq. (6) is then expressed in matrix form as:
@@
It is important to note that the structure of the matrix least squares problem for the general parallel RC is similar to Eq. (8). However, in the general case, we solve for each $W_{o}^{(k)}$ independently, considering the corresponding $\mathbf{X}^{(k)}$ and $\mathbf{\hat{Y}}^{(k)}$, rather than stacking them as $\mathbf{X}$ and $\mathbf{\hat{Y}}$.
@@ 2.3 Testing Process and Quality Metric
Two primary prediction modes are used to assess the predictive capability of RC [46]. The first mode is ”teacher forcing” or one-step prediction, where the reservoir is fed with test input data $\mathbf{u}_{+}^{(k)}=\begin{pmatrix}u_{0}^{(k)}&u_{1}^{(k)}&\ldots\end{pmatrix}$. Note that the testing input data are ”ground-truth” data derived from the same partial differential equation as the training data, but with different times or initial conditions. The hidden states evolve according to Eqs. (1) and (3). The one-step prediction is obtained by applying Eqs. (2) and (4) using the learned output matrix $W_{o}$. It is important to note that, in this mode, the RC prediction $\hat{y}_{j}^{(k)}$ depends on the reservoir states $x_{j}^{(k)}$, but is independent of the previous time step’s prediction, $\hat{y}_{j-1}^{(k)}$. This mode is valuable for assessing the model’s generalization ability to unseen inputs.
The second prediction mode is the generative or autonomous mode. In this mode, the reservoir is initially fed with ground-truth input data $\mathbf{u}_{+}^{(k)}$ until a specific time index $j_{0}$. Subsequently, the RC prediction at time $j_{0}$, $\hat{y}_{j_{0}}^{(k)}$, is used as the input states $\hat{u}_{j_{0}}^{(k)}$ to update the hidden states $x_{j_{0}+1}^{(k)}$ and generate the prediction $\hat{y}_{j_{0}+1}^{(k)}$. This iterative process continues until the desired prediction steps are achieved. For our parallel RC, there is a subtle difference between the RC prediction $\hat{y}_{j}^{(k)}$ and the feedback input $\hat{u}_{j}^{(k)}$ due to the overlapping grid. Specifically, the input $\hat{u}_{j_{0}}^{(k)}$ should be assembled using $\hat{y}_{j_{0}}^{(k)}$ and its adjacent predictions $\hat{y}_{j_{0}}^{(k\pm 1)}$, as illustrated in Fig. 1. It is important to note that in the generative mode, the error will propagate during subsequent predictions, making this prediction more challenging than the teacher-forcing mode. Although the generative mode is commonly used to assess RC performance, we aim to understand both types of prediction modes.
@@
3 RC Prediction for Higher Order Rogue Waves
In this section, we present the results of applying the parallel RC model to predict the occurrence of rogue waves. We begin by explaining how we generate the training data. Then, we outline the implementation of our training procedure and discuss the results for both one-step and generative predictions.
@@ 3.1 Generation of Ground Truth Data
A well-known analytic solution of the NLS equation is the Akhmadiev breather [47], given by:
$\psi(\xi,t)=\left[1+\frac{2(1-2a)\cosh(\gamma t)+i\gamma\sinh(\gamma t)}{\sqrt{2a}\cos(\Omega\xi)-\cosh(\gamma t)}\right]e^{it},$
@@
where $\Omega$ is the fundamental frequency corresponding to $a\approx 0.4802$ or $\Omega\approx 0.39799$. With this choice, there are five unstable modes corresponding to frequencies $\Omega,2\Omega,\dots,5\Omega$, with respective growth factors $(\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5})=(0.39,0.73,0.958,0.964,0.199)$. We set $A_{1}=10^{-4}$, while $A_{0}=\sqrt{1-2A_{1}^{2}}$. The simulation uses a fixed timestep of $\Delta t=2\times 10^{-4}$.
The length of the spatial domain is one spatial period, $[-L/2,L/2]$, with the primary wavelength $L=2\pi/\Omega$. The spatial domain consists of 256 spatial nodes, with periodic boundary conditions on the lateral sides. After the first formation of the Akhmadiev breather, higher-order recurrences are created due to higher-order dynamics modes located in the modulation instability range [51, 52, 53]. Thus, a complex recurrence pattern grows and collapses as modes with different growth factors and phases interact nonlinearly.
3.2 Training and One-Step Prediction
@@
For our RC experiment, we separate the real and imaginary parts of the NLS, yielding 512 spatial nodes. We sample the data such that $\Delta t_{RC}=5\times 10^{-3}$ time units, equivalent to taking one sample every 25 time steps. Initially represented in quadruple precision, the simulated data is truncated to double precision. To eliminate the initial build-up of the first breather, we exclude 15 time units from the initial simulation. Subsequently, the next 70 time units of the data are used for training, while the following 35 time units serve as testing data. The division between training and testing data is shown in Fig. 2, with the white line indicating the boundary. We refer to this boundary as $t=0$ in the following results.
@@
For the training process, we set the regularization strength $\beta=10^{-4}$ and use $100$ washout steps ($N_{w}=100$). Remarkably, these hyperparameters are not highly sensitive. We introduce white Gaussian noise with a variance of $\sigma=0.02$ only to the system input data $\mathbf{u}^{(k)}$. We observed an improvement in the autonomous prediction when noise was introduced, although the optimal noise value remains unclear [22]. Consistent with the findings in [45], we observe that the translation invariance procedure significantly improved the autonomous prediction in our case. We will refer to this RC model as RC_AB.
After training, we generate a teacher-forced reservoir by feeding the test data, $u_{j}^{(k)}$, for all $k$ and $j\geq 0$. Based on these reservoir states, we compute a one-step prediction at each time step $\hat{y}_{j+1}$ and compare it with the real test data $y_{j+1}$. In this mode, the one-step error at a given time instance does not propagate to the next prediction. The NRMSE for both training and one-step predictions are shown in Fig. 3, along with the maximum wave envelope for comparison with Fig. 2. The figure shows that the overall error in the training procedure is higher than in the one-step prediction, which can be attributed to high noise in the training data. The testing data also show that the error overshoots during the modulation instability, highlighting the difficulty of predicting nonlinear instability. Nevertheless, the error remains small, and an accurate one-step prediction is achieved even during the instability.
Figure 3: Trained RC has low NRMSE for one-step forecasting. The top and bottom panels show the time trace of training NRMSE and NRMSE of one-step testing predictions, respectively. The black line represents the NRMSE, while the red line represents the maximum wave envelope.
@@ 3.3 Autonomous Prediction
We investigate the autonomous prediction performance of the RC on the test data. In this analysis, we vary the initial time of our autonomous prediction, $t_{0}$, and observe the prediction horizon. As noted previously, in each autonomous prediction, we always feed the RC with ground truth data before $t_{0}$ (see Section 2.3). The results show that the reservoir has a long prediction horizon, as illustrated in Fig. 4. However, the RC has difficulty forecasting near the peak of the modulation instability. Nevertheless, the RC can detect rogue waves approximately two time units before their occurrence.
@@
As an alternative, we consider the maximum modulation instability growth factor, $\gamma_{\text{max}}=\gamma_{4}$, as the characteristic timescale. This choice is motivated by two main reasons: first, this maximum growth rate is equivalent to the largest Lyapunov exponent in the case of a plane wave solution, and second, the maximum exponential growth of a rogue wave follows this rate, analogous to the exponential divergence of trajectories in chaotic systems. We support this claim by calculating the local largest Lyapunov exponent, as detailed in B . Given that $\gamma_{\text{max}}=0.964$, the timescale shown in Fig. 4 is nearly aligned with this characteristic timescale.
Figure 4: The prediction horizon (PH) of autonomous prediction varies at different start times $t_{0}$. Close to the peak of modulation instability, PH becomes shorter.
@@
Figure 5: Autonomous RC_AB simulation for offset time instances. Actual data are plotted for comparison. Note how the amplitude difference diverges. Nevertheless, RC_AB can forecast the dynamics. The autonomous prediction starts at $t_{0}=13.4$. The green curves are from RC_Cm, see Section 5.
Figure 6: Kinetic Energy (KE), Potential Energy (PE), and Hamiltonian (Ham) of the autonomous RC_AB, RC_Cm (see Section 5), and the NLS simulation. The learned RCs behave like a dissipative system. The peak of the kinetic energy (around $t=15.4$) marks the breather occurrence.

### v2

@@ 1 Introduction
Reservoir Computing (RC) is a specific variant of the recurrent neural network (RNN) in machine learning [22]. RC can be perceived as a three-layer RNN: an input layer, an internal layer state (a reservoir), and an output layer. In contrast to the conventional RNN, the reservoir parameters are fixed, sparse, and random. Surprisingly, despite having a simplified learning procedure compared to RNNs, RC has been reported to be more powerful [23, 24, 25]. The framework has successfully modeled both short-term and long-term dynamics of chaotic systems and captured its attractor [26, 27, 28, 29, 30]. An extension of RC for learning spatiotemporal problems was presented in [27, 25, 31, 32, 33, 34, 35]. Recent studies [36, 37] have introduced knowledge-based models within the RC framework. These advances highlight the versatility and potential of RC in addressing a wide range of spatiotemporal challenges. Another motivation for RC lies in its potential for physical implementation, as reviewed in [38, 39, 40], which has the potential to achieve both computational speedup and improved energy efficiency.
This research investigates the prediction and forecasting capabilities of RC to predict short-term nonlinear dynamics of integrable systems modeled by the NLS equation. Despite various studies on RC in chaotic spatiotemporal systems, little attention has been given to its application to Hamiltonian counterparts. An investigation into RC’s ability to replicate the phase plane of Hamiltonian systems is reported in [41], although this study is limited to low-dimensional systems. Only [32] attempted to simulate NLS using RC, focusing on periodic Akhmediev breathers, Kutnezov–Ma solitons, and second-order periodic breather collisions. Despite this insight, the similarity between the training and testing data in [32] raises questions about the model’s generalizability.
@@
The primary objective of this research is to advance the understanding of RC’s forecasting capabilities for non-periodic waves, particularly rogue waves. A key focus lies in evaluating its performance using testing data from a different initial condition, aiming to thoroughly assess its robustness and applicability in varying scenarios. While several studies have demonstrated RC’s ability to predict unseen dynamics in chaotic dissipative systems [42, 29, 43], its performance in other systems, particularly Hamiltonian systems,remains largely unexplored due to the absence of an attractor. Our main contribution to the project is the revelation that RC can provide reasonable predictions, even when operating within different phase spaces in the integrable spatiotemporal system. Furthermore, we introduce a modified forecasting approach that incorporates real data assimilation, significantly improving RC’s predictive performance. This result paves the way for future research to better understand and improve the long-term predictive capabilities of RC within Hamiltonian systems.
This report is organized into the following sections. Section 2 presents a brief introduction to parallel RC and details the training and testing procedures used in the study. In Section 3, we provide a detailed explanation of the training data and the prediction results. The main focus of our study is explained in Section 4 and Section 5. In these sections, we employ different testing data to assess the generalizability of the RC model trained in the previous section for higher-order breathers and random ocean waves. Section 6 discusses procedures to extend the prediction horizon in autonomous prediction. We conclude in Section 7 and provide an overview of the future research direction.
2 Reservoir Computing Framework
@@ 2.1 Parallel Reservoir Computing
We will refer to the pair $(f,g)$ as an RC, where the input states are denoted as $u_{j}\in\mathbb{R}^{d_{I}}$ and the hidden states as $x_{j}\in\mathbb{R}^{d_{S}}$. As evident from Eqs. (1)–(2), RC is driven by a sequence of input states $\mathbf{u}=(\ldots,u_{-1},u_{0})$, complemented by an initial hidden state $x_{-\infty}$. We denote $\mathbf{x}=(\ldots,x_{-1},x_{0})$ and write $\mathbf{x}=f(\mathbf{u},x_{-\infty})$ if each element of the sequence $\mathbf{x}$ is obtained by applying the operation in (1) to the input sequence $\mathbf{u}$. Consequently, the output sequence is written as $\mathbf{y}=(f,g)(\mathbf{u})$.
When the effect of the initial hidden state $x_{-\infty}$ becomes negligible after a certain number of iterations, implying that the hidden state depends solely on the input sequence, the dynamical system exhibits the Echo State Property (ESP) [23]. Technically, if $\mathbf{x}=f(\mathbf{u},x_{-\infty})$ and $\tilde{\mathbf{x}}=f(\mathbf{u},\tilde{x}_{-\infty})$ are two sequences of hidden states corresponding to different initial conditions, the ESP condition implies that $x_{0}=\tilde{x}_{0}$.
@@
The choice of RC, as described by Eqs. (3)–(4), takes into account the symmetry inherent in the underlying data to be simulated. Specifically, the RC exhibits symmetry such that $(f,g)(\mathbf{u})=-(f,g)(-\mathbf{u})$. The importance of preserving the symmetry of the RC output, which mirrors the symmetry of the original system, has been emphasized in previous works [27, 45]. Acknowledging and preserving this symmetry is crucial for ensuring the fidelity of the simulated dynamics and improving the ability of the RC model to accurately capture the underlying patterns and behaviors of the system being studied.
@@ 2.2 Training Procedure
Figure 1: Illustration of the parallel RCs architecture. The spatial grid is divided into overlapping grids $u^{(k)}$ as input to RCs and non-overlapping grids $y^{(k)}$ as the output.
@@
where $N_{w}$ is the number of initialization steps used to ”warm up” the RCs. The first $N_{w}$ hidden states are discarded to eliminate any dependencies on initialization. The terms $\hat{y}_{j}^{(k)}$ represent the true output data, and $\beta$ is the Tikhonov parameter that controls the regularization strength of the least-square problem. The norm of the matrix $W_{o}$ is computed using the Frobenius norm.
@@
Let us define $\mathbf{X}^{(k)}$ and $\mathbf{\hat{Y}}^{(k)}$ as matrices whose columns are the vectors $x^{(k)}_{j}$ and $\hat{y}^{(k)}_{j}$, respectively, for $-N+N_{w}\leq j\leq 0$. Furthermore, we construct the matrix $\mathbf{X}$ by stacking the matrices $\mathbf{X}^{(k)}$ column-wise, and similarly for $\hat{\mathbf{Y}}$. The loss function (6) is then expressed in matrix form as:
@@
It is important to note that the structure of the matrix least squares problem for the general parallel RC is similar to Eq. (8). However, in the general case, we solve for each $W_{o}^{(k)}$ independently, considering the corresponding $\mathbf{X}^{(k)}$ and $\hat{\mathbf{Y}}^{(k)}$, rather than stacking them as $\mathbf{X}$ and $\hat{\mathbf{Y}}$.
@@ 2.3 Testing Process and Quality Metric
Two primary prediction modes are used to assess the predictive capability of RC [46]. The first mode is ”teacher forcing” or one-step prediction, where the reservoir is fed with test input data $\mathbf{u}_{+}^{(k)}=\begin{pmatrix}u_{0}^{(k)}&u_{1}^{(k)}&\ldots\end{pmatrix}$. Note that the testing input data are ”ground-truth” data derived from the same partial differential equation as the training data, but with different times or initial conditions. The hidden states evolve according to Eqs. (1) and (3). The one-step prediction is obtained by applying Eqs. (2) and (4) using the learned output matrix $W_{o}$. It is important to note that, in this mode, the RC prediction $\hat{y}^{(k)}_{j}$ depends on the reservoir states $x^{(k)}_{j-1}$ but is independent of the previous time step’s prediction, $\hat{y}^{(k)}_{j-1}$. This mode is valuable for assessing the model’s generalization ability to unseen inputs.
The second prediction mode is the generative or autonomous mode. In this mode, the reservoir is initially fed with ground-truth input data $\mathbf{u}_{+}^{(k)}$ until a specific time index $j_{0}$. Subsequently, the RC prediction at time $j_{0}$, $\hat{y}_{j_{0}}^{(k)}$, is used as the input states $\hat{u}_{j_{0}}^{(k)}$ to update the hidden states $x_{j_{0}+1}^{(k)}$ and generate the prediction $\hat{y}_{j_{0}+1}^{(k)}$. This iterative process continues until the desired prediction steps are achieved. For our parallel RC, there is a subtle difference between the RC prediction $\hat{y}_{j}^{(k)}$ and the feedback input $\hat{u}_{j_{0}}^{(k)}$, due to the overlapping grid. Specifically, the input $\hat{u}_{j_{0}}^{(k)}$ should be assembled using $\hat{y}_{j_{0}}^{(k)}$ and its adjacent predictions $\hat{y}_{j_{0}}^{(k\pm 1)}$, as illustrated in Fig. 1. It is important to note that in the generative mode, the error will propagate during subsequent predictions, making this prediction more challenging than the teacher-forcing mode. Although the generative mode is commonly used to assess RC performance, we aim to understand both types of prediction modes.
@@
3 RC Prediction for Higher-Order Rogue Wave
In this section, we present the results of applying the parallel RC model to predict the occurrence of rogue waves. We begin by explaining how we generate the training data. Then, we outline the implementation of our training procedure and discuss the results for both one-step and generative prediction.
@@ 3.1 Generation of Ground Truth Data
A well-known analytic solution of the NLS equation is the Akhmediev breather [47], given by:
$\psi(\xi,t)=\left[1+\frac{2(1-2a)\cosh\left(\gamma t\right)+i\gamma\sinh\left(\gamma t\right)}{\sqrt{2a}\cos\left(\Omega\xi\right)-\cosh\left(\gamma t\right)}\right]e^{it},$
@@
where $\Omega$ is the fundamental frequency corresponding to $a\approx 0.4802$ or $\Omega\approx 0.39799$. With this choice, there are five unstable modes corresponding to frequencies $\Omega,2\Omega,\dots,5\Omega$, with respective growth factors $(\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5})=(0.39,0.73,0.958,0.964,0.199)$. We set $A_{1}=10^{-4}$, while $A_{0}=\sqrt{1-2A_{1}^{2}}$. The simulation uses a fixed timestep $\Delta t=2\cdot 10^{-4}$.
The length of the spatial domain is one spatial period, $[-L/2,L/2]$, with the primary wavelength $L=2\pi/\Omega$. The spatial domain consists of 256 spatial nodes, with periodic boundary conditions on the lateral sides. After the first formation of the Akhmediev breather, higher-order recurrences are created due to higher-order dynamics modes located in the modulation instability range [51, 52, 53]. Thus, a complex recurrence pattern grows and collapses as modes with different growth factors and phases interact nonlinearly.
3.2 Training and One Step Prediction
@@
For our RC experiment, we separate the real and imaginary parts of the NLS, yielding 512 spatial nodes. We sample the data such that $\Delta t_{RC}=5\times 10^{-3}$ time units, equivalent to taking one sample every 25 time steps. Initially represented in quadruple precision, the simulated data is truncated to double precision. To eliminate the initial build-up of the first breather, we exclude 15 time units from the initial simulation. Subsequently, the next 70 time units of the data are used for training data, while the following 35 time units serve as testing data. The division between training and testing data is shown in Fig. 2, with the white line indicating the boundary. We refer to this boundary as $t=0$ in the following results.
Figure 3: Trained RC has low NRMSE for one-step forecasting. The top and bottom panels show the time trace of training NRMSE and NRMSE of one-step testing prediction respectively. The black line represents the NRMSE, while the red line is the maximum wave envelope
@@
For the training process, we set the regularization strength $\beta=10^{-4}$ and use $100$ washout steps ($N_{w}=100$). Remarkably, these hyperparameters are not highly sensitive. We introduce white Gaussian noise with a variance of $\sigma=0.02$ only to the system input data $\mathbf{u}^{(k)}$. We observed an improvement in the autonomous prediction when noise was introduced, although the optimal noise value remains unclear [22]. Consistent with the finding in [45], we observe that the translation invariance procedure significantly improved the autonomous prediction in our case. We will refer to this RC model as RC_AB.
After training, we generate a teacher-forced reservoir by feeding the test data, $u_{j}^{(k)}$, for all $k$ and $j\geq 0$. Based on these reservoir states, we compute a one-step prediction at each time step $\hat{y}_{j+1}$ and compare it with the real test data $y_{j+1}$. In this mode, the one-step error at a given time instance does not propagate to the next prediction. The NRMSE for both training and one-step prediction are shown in Fig. 3, along with the maximum wave envelope for comparison with Fig. 2. The figure shows that the overall error in the training procedure is higher than in the one-step prediction, which can be attributed to high noise in the training data. The testing data also show that the error overshoots during the modulation instability, highlighting the difficulty of predicting nonlinear instability. Nevertheless, the error remains small, and an accurate one-step prediction is achieved even during the instability.
@@ 3.3 Autonomous Prediction
We investigate the autonomous prediction performance of the RC on the test data. In this analysis, we vary the initial time of our autonomous prediction, $t_{0}$, and observe the prediction horizon. As noted previously, in each autonomous prediction, we always feed the RC with ground truth data before $t_{0}$ (see Sec. 2.3). The results show that the reservoir has a long prediction horizon, as illustrated in Fig. 4. However, the RC has difficulty forecasting near the peak of the modulation instability. Nevertheless, the RC can detect rogue waves approximately two time units before their occurrence.
@@
As an alternative, we consider the maximum modulation instability growth factor, $\gamma_{\text{max}}=\gamma_{4}$, as the characteristic timescale. This choice is motivated by two main reasons: first, this maximum growth rate is equivalent to the largest Lyapunov exponent in the case of a plane wave solution, and second, the maximum exponential growth of a rogue wave follows this rate, analogous to the exponential divergence of trajectories in chaotic systems. We support this claim by calculating the local largest Lyapunov exponent, as detailed in Appendix B . Given that $\gamma_{\text{max}}=0.964$, the timescale shown in Fig. 4 is nearly aligned with this characteristic timescale.
Figure 4: The prediction horizon (PH) of autonomous prediction varies at different start times $t_{0}$. Close to the peak of modulation instability, PH becomes short.
@@
Figure 5: Autonomous RC_AB simulation for offset time instances. Actual data are plotted for comparison. Note how the amplitude difference diverges. Nevertheless, RC_AB can forecast the dynamics. The autonomous prediction starts at $t_{0}=13.4$. The green curves are from RC_Cm, see Sec. 5.
Figure 6: Kinetic Energy (KE), Potential Energy (PE), and Hamiltonian (Ham) of the autonomous RC_AB, RC_Cm (see Sec. 5), and the NLS simulation. The learned RCs behave like a dissipative system. The peak of the kinetic energy (around $t=15.4$) marks the breather occurrence.